# Optical Physiological Signals

### v1

@@ D.2 Other Evaluation Metrics
Appendix E Effect of Skin Tone
@@
Appendix F Extended related work
@@
However, despite the popularity of SSL, there are no widely used pre-trained models for PPG data. A recent study (Abbaspourazad et al., 2023) demonstrated that embeddings derived from ECG and PPG signals can generalize across multiple health outcomes using proprietary large-scale data. While this work showcased the potential of foundation models for physiological signals, it was based on a single dataset and device (Apple Watch) while the models were not released, limiting its practical use in the research community. Similarly, REGLE’s work (Yun et al., 2024) on the UK Biobank dataset showed that embedding PPG signals can improve genetic discovery and risk prediction outcomes. Although parts of that model and pipeline are public, the dataset is not openly accessible, and the primary goal was not to create a foundation model for PPG but rather to focus on genetics. Another work on the same data showed that PPG embeddings are promising for cardiovascular risk prediction (Weng et al., 2024). SiamQuality (Ding et al., 2024) also trained an unreleased model on 36 million PPG signals using proprietary data. Importantly, most of these works pre-trained on a single-device dataset and did not explore out-of-domain datasets or conduct transfer learning experiments, which are crucial for assessing the true generalizability of foundation models. These studies highlight the potential of PPG-based foundation models but also underscore the need for openly available, pre-trained models that can be widely used and adapted by the research community.
@@
In contrast to these generic approaches, our work takes a domain-specific focus. We curate a large pre-training and evaluation benchmark dedicated exclusively to PPG data. While knowledge gained from generic time series foundation models may transfer to domain-specific tasks like PPG, we expect the performance to be limited compared to a model trained specifically on PPG data. Furthermore, foundation models for ECG (McKeen et al., 2024; Song et al., 2024) or EEG (Yuan et al., 2024b) have shown promise but transferring from one domain-specific model (e.g., ECG) to another (PPG) is likely to be even more challenging, as the underlying signal characteristics can be quite different. Building on the increasing interest in modality-specific foundation models, our specialized approach allows us to capture nuances and complexities specific to PPG signals.
@@
Appendix G Extended Discussion
Our results show that PaPaGei outperforms baselines in at least 14 out of 20 tasks, with average classification and regression improvements of 4.7%-6.3% and 2.9%-4.9%, respectively (Section 5.1). PaPaGei-S performed best for cardiovascular parameters like BP, Hypertension, and Avg. HR, which are closely linked to components such as sVRI and IPA (Liang et al., 2018b). Additionally, PaPaGei-P surpassed pre-trained baselines like Moment and is well-suited for tasks such as Smoking and Arousal.

### v2

@@ D.3 Ablation Results
Table 14: PaPaGei component ablation study results.
sVRI sVRI + SQI sVRI + IPA Full
Classification - AUROC ($\uparrow$)
ICU Admission 0.79 0.75 0.78 0.79
Mortality 0.67 0.65 0.67 0.67
Smoker 0.59 0.61 0.60 0.61
Pregnancy stage 0.78 0.73 0.72 0.78
Hypertension 0.77 0.72 0.75 0.77
Sleep Disordered Breathing 0.62 0.53 0.64 0.70
Mood Disturbance 0.53 0.56 0.55 0.56
Valence 0.54 0.55 0.53 0.56
Arousal 0.44 0.51 0.49 0.55
Regression - MAE ($\downarrow$)
Apnea/Hypopnea Index $>$ 3% 13.36 13.74 13.42 12.97
Apnea/Hypopnea Index $>$ 4% 11.01 11.43 11.29 10.56
Gestation Age 6.18 6.32 6.15 6.05
Systolic BP (VV) 14.62 15.97 15.33 14.65
Diastolic BP (VV) 8.32 8.76 9.04 8.29
Systolic BP (PPG-BP) 15.03 14.39 16.15 14.39
Diastolic BP (PPG-BP) 9.12 8.76 9.06 8.71
Average HR 4.00 5.88 4.26 4.00
HR 11.51 11.97 11.88 11.53
D.4 Statistical Significance of Model Comparison
In addition to confidence intervals, we perform the following steps to evaluate the significance across models on a per task basis (Tables 3 & 4). First, we ran the Friedmann Chi Square test, and identified statistically significant differences across PaPaGei and the baseline models at $p<0.05$. Next, we created critical difference (CD) diagrams to rank the best performing models, as suggested by the literature to compare models over multiple datasets [^6] https://scikit-posthocs.readthedocs.io/en/latest/tutorial.html#critical-difference-diagrams (Demšar, 2006). The CDs indicate that PaPaGei performs the best across both classification and regression tasks. Furthermore, it has a statistically significant average rank as indicated by the lack of horizontal line.
From the critical difference diagrams we observe that PaPaGei is significantly better across classification (Figure 22) and regression (Figure 23) tasks. This arises because PaPaGei is the highest ranking model across most tasks. Furthermore, we observe Moment is a strong model across both classification and regression tasks. Whereas Chronos and TF-C perform well for classification tasks only.
Figure 22: Critical Difference Diagram for Classification Tasks. The axis represents the average rank of the model. The horizontal connector lines indicate no significant differences between the models.
Figure 23: Critical Difference Diagram for Regression Tasks. The axis represents the average rank of the model. The horizontal connector lines indicate no significant differences between the models.
We conduct additional statistical significance comparisons using a structured approach. First, we randomly sample a score from within the confidence intervals for each task across all models. Next, we apply the CD ranking procedure to the sampled scores. This process is repeated 1,000 times, and the ranks are averaged. The entire experiment is conducted five times for Tables 3 and 4, with the results presented in Figure 24. The colored cells indicate that PaPaGei is statistically significant compared to the respective model at $p<0.05$. Our findings show that PaPaGei consistently achieves the best average rank, ranging between 0.82-0.90 for AUROC and 0.19-0.25 for MAE. Across 35 comparisons (PaPaGei vs. the other models, repeated five times), PaPaGei demonstrates significant improvements in 30 out of 35 AUROC comparisons and 32 out of 35 MAE comparisons. Among the baseline models, we acknowledge that Chronos and TF-C are strong competitors capable of performing comparably to PaPaGei.
Figure 24: Bootstrap ranking repeated for five experiments: AUROC (left) and MAE (right). Colored cells indicate that PaPaGei is significant to the baseline at $p<0.05$.
D.5 Statistical assessment between IPA and SQI
Recall that we incorporate SQI to handle situations where the dicrotic notch cannot be computed due to poor-signal quality or different morphologies. We performed a permutation test (Rice, 2008) to statistically evaluate PPG segments where IPA is unavailable. By splitting SQI values into no IPA and IPA groups and testing significance over 1000 permutations, we observed statistically significant differences ($p<0.05$) in both mean (+0.18) and median (+0.32) SQI values, with the IPA group having larger SQI. These findings empirically motivate SQI’s ability to handle limited PPG morphology.
Appendix E Additional Baselines: Demographics & PPG Morphology Features
In this section, we evaluate the effectiveness of demographics (Demo: age, sex) and PPG morphology (sVRI, IPA, SQI) to predict both regression (ridge) and classification (logistic regression) tasks: (1) Ablation Study: We compared PaPaGei with three baselines—demographics alone, PPG features alone, and demographics + PPG features. Our results show that while demographics is a stronger baseline than statistical features, PaPaGei outperforms the demographics + PPG baseline in 14 out of 18 tasks. (2) Effect of Demographics: We trained a model combining PaPaGei-S with demographics. The results indicate that incorporating demographic features with PaPaGei-S creates a stronger model than using PaPaGei-S alone.
Table 15: Demographics & PPG Morphology Baseline Results.
Stat. Features Demo PPG Demo + PPG PaPaGei-S or -P PaPaGei-S + Demo
Classification - AUROC ($\uparrow$)
ICU Admission 0.71 [0.65-0.78] 0.64 [0.60-0.68] 0.59 [0.54-0.64] 0.66 [0.61-0.70] 0.79 [0.75-0.82] 0.77 [0.74-0.81]
Mortality 0.57 [0.54-0.61] 0.66 [0.61-0.69] 0.57 [0.53-0.61] 0.66 [0.61-0.69] 0.67 [0.63-0.70] 0.70 [0.67-0.74]
Smoker 0.63 [0.58-0.67] 0.64 [0.59-0.70] 0.56 [0.52-0.62] 0.66 [0.61-0.71] 0.64 [0.58-0.69] 0.62 [0.56-0.68]
Pregnancy stage 0.64 [0.62-0.67] 0.52 [0.50-0.55] 0.55 [0.53-0.57] 0.56 [0.53-0.58] 0.78 [0.75-0.80] 0.78 [0.76-0.80]
Hypertension 0.66 [0.47-0.83] 0.77 [0.65-0.88] 0.53 [0.40-0.68] 0.77 [0.65-0.88] 0.77 [0.68-0.87] 0.80 [0.70-0.89]
SDB 0.32 [0.14-0.55] 0.46 [0.31-0.62] 0.70 [0.57-0.84]
Mood Disturbance 0.54 [0.31-0.77] 0.54 [0.30-0.80] 0.64 [0.42-0.85] 0.63 [0.36-0.87] 0.56 [0.33-0.77] 0.52 [0.25-0.78]
Valence 0.52 [0.49-0.55] 0.57 [0.54-0.60] 0.44 [0.41-0.47] 0.57 [0.54-0.60] 0.56 [0.54-0.59] 0.55 [0.53-0.58]
Arousal 0.55 [0.53-0.58] 0.54 [0.52-0.58] 0.51 [0.48-0.54] 0.54 [0.52-0.58] 0.58 [0.55-0.61] 0.58 [0.54-0.59]
Regression - MAE ($\downarrow$)
Apnea/Hypopnea Index $>$ 3% 15.31 [13.63-17.14] 14.53 [13.29-15.84] 15.09 [14.01-16.54] 14.40 [13.12-15.61] 12.97 [11.87-14.05] 12.35 [11.27-13.46]
Apnea/Hypopnea Index $>$ 4% 12.52 [10.92-14.14] 12.28 [11.19-13.39] 12.65 [11.57-13.83] 12.17 [11.10-13.39] 10.56 [9.59-11.62] 10.47 [9.53-11.50]
Gestation Age 7.15 [6.99-7.34] 7.69 [7.61-7.77] 7.61 [7.51-7.70] 7.59 [7.51-7.68] 6.05 [5.91-6.17] 6.02 [5.88-6.17]
Systolic BP (VV) 15.76 [13.67-18.36] 14.96 [13.21-17.35] 15.82 [13.48-18.31] 15.01 [13.30-17.86] 14.65 [12.50-16.78] 14.27 [11.92-16.44]
Diastolic BP (VV) 9.75 [7.16-11.27] 8.75 [6.48-9.77] 9.20 [7.21-10.71] 8.78 [7.10-10.25] 8.29 [6.61-10.22] 8.26 [6.64-10.16]
Systolic BP (PPG-BP) 15.50 [11.68-20.25] 13.71 [11.33-15.95] 15.76 [13.36-18.30] 13.74 [11.37-16.09] 13.60 [10.65-16.51] 13.20 [11.47-15.66]
Diastolic BP (PPG-BP) 9.35 [7.44-11.66] 9.26 [7.89-10.68] 9.36 [7.95-10.92] 9.28 [8.00-10.56] 8.71 [7.18-10.01] 8.61 [7.34-9.88]
Average HR 7.01 [5.48-8.89] 9.12 [7.86-10.61] 8.07 [6.60-9.71] 8.23 [6.82-9.78] 3.47 [2.74-4.32] 4.00 [3.35-4.73]
HR 13.07 [12.90-13.23] 15.18 [15.03-15.33] 16.75 [16.60-16.90] 14.46 [14.32-14.62] 10.92 [10.80-11.04] 12.38 [11.90-12.96]
From Table 15, we observe the following classification performance (Positive is better): ICU (+0.13), Mortality (+0.01), Smoker (-0.02), Pregnancy Stage (+0.22), Hypertension (0.00), SDB (no demographics), Mood Disturbance (-0.08), Valence (-0.01), Arousal (+0.04). Regression Tasks (Negative is better): AHI $>$ 3% (-1.43), AHI $>$ 4% (-1.61), gestation age (-1.54), SBP-VV (-0.31), DBP-VV (-0.46), SBP (-0.11) , DBP (-0.65), Avg. HR (-4.07), HR (-2.93). PaPaGei-S performs better for real-time sleep and cardiovascular outcomes such as sleep apnea, heart rate and blood pressure, respectively. In particular, we notice that outcomes such as heart rate benefit substantially from PPG rather than demographics. Demographics are useful in tasks without real-time dependence such as smoking, which is established to be associated with age and sex (Chung et al., 2020). Importantly, demographics do not add much to already homogeneous populations. For example, consider the NuMoM2B dataset which has women within a specific age range. Here, we observe that PaPaGei obtains much higher AUROC and MAE than the supervised baselines.
Furthermore, We observe that adding demographics to PaPaGei-S embeddings improves over PaPaGei in the following tasks: Mortality (+0.03), Hypertension (+0.03), AHI $>$ 3% (-0.62), AHI $>$ 4% (-0.09), gestation age (-0.03), SBP VV (-0.38), DBP VV (-0.03), SBP (0.40), DBP (0.10). Based on these results, PaPaGei-S + Demo is a stronger model in many cases. Importantly, these results indicate that PaPaGei-S embeddings learn features that are complementary to demographics are not simply proxies for age or sex. However, it is important to note that while demographic features can be valuable for personalization, they may not always be readily available, and in reality, we cannot use them in isolation to predict real-time outcomes such as blood pressure or heart rate. Therefore, our PaPaGei models are designed to function effectively with real-time sensor data alone, ensuring their applicability in situations where complete demographic information is not accessible.
These findings underscore an important point: demographic features are not competing with PaPaGei but rather complement it, as previously established in studies including demographics with sensor data (Spathis et al., 2022). This highlights the synergistic potential of combining PaPaGei’s advanced feature extraction with demographic context for improved task performance.
Predicting Demographics Targets. Using the PaPaGei features, we predict downstream demographics such as age and sex (Table 16). PaPaGei-S achieves 7.78 MAE in age regression, 0.85 accuracy in age classification, and 0.79 accuracy in sex classification. Although our results trail larger closed studies (Abbaspourazad et al., 2023) by 2.18, 0.05, and 0.13 for segment-level SSL, and by 5.59, 0.12, and 0.25 for patient-level SSL, they mark an advancement in open-source efforts. The superior performance of Abbaspourazad et al. (2023) can be attributed to two primary factors. First, the patient-level positive pair strategy achieves the best performance across all tasks. This approach encourages the model to form distinct clusters for each patient, effectively capturing demographic factors such as age and sex. In contrast, a segment-level approach pushes the model to cluster similar segments across individuals with varying demographics, potentially mixing demographic-specific information. Second, the single-device setup with a larger dataset is useful for effective model training (Table 17). Conversely, our evaluation, which spans three devices and utilizes smaller datasets, must handle greater data heterogeneity, thus making it more challenging.
Table 16: Predicting personal characteristics with embeddings. Downstream prediction on age regression, age classification, and sex classification in our pre-training datasets (VitalDB, MESA, MIMIC-III). The regression and classification tasks are reported using MAE and AUROC, respectively. Note: training and testing are conducted with completely different cohorts in the two studies, hence comparisons are difficult.
Study Age Regression ($\downarrow$) Age Classification ($\uparrow$) Sex Classification ($\uparrow$)
Abbaspourazad et al. (2023)(Patient) 3.19 0.97 0.99
Abbaspourazad et al. (2023)(Segment) 6.60 0.90 0.87
PaPaGei-S (Ours) 8.78 [8.47-8.09] 0.85 [0.83-0.87] 0.74 [0.72-0.76]
Appendix F Additional Prediction Plots
The regression plots to evaluate the agreement between true and predicted values in shown in Figure 25. From the Figure, we observe that PaPaGei’s predictions are more aligned to the true values for AHI $>$ 3% ($R^{2}=0.28$), Avg. HR ($R^{2}=0.79$), gestation age ($R^{2}=0.28$), SBP ($R^{2}=0.36$), and DBP ($R^{2}=0.22$). Moreover, from the distribution plots in Figure 25, we notice that PaPaGei has stronger overlap for AHI $>$ 4%, Avg. HR and DBP, indicating its ability to capture the tails for these tasks. Interestingly, we notice that all models are unable to capture the bi-modal nature of the gestation age measurements. Here, Chronos performs better than other methods to capture readings from the first visit.
((a)) AHI $>$ 4% predictions
((b)) Avg. HR predictions
((c)) Gestation Age
((d)) Systolic BP (PPG-BP)
((e)) Diastolic BP (PPG-BP)
Figure 25: Regression plots and prediction distribution of different models compared to ground truth for (a) Apnea/Hypopnea Index $>$ 4%, (b) Average Heart Rate, (c) Gestation Age, (d) Systolic BP (PPG-BP), and (e) Diastolic BP (PPG-BP). $R^{2}$ is the coefficient of determination and $m$ is the correlation slope.
Appendix G Effect of Skin Tone
We present the skin tone analysis in a more granular way in Figure 26. Here, PaPaGei-S clearly performs better than PaPaGei-P in most cases. Overall, we notice that PaPaGei-S is good for lighter skin tones in the 1-2 range for SBP and 2-3 range for DBP. While PaPaGei-S does not perform the best for darker skin tones, it’s performance is comparable to other models for skin tone ratings of 4 and 5. Overall, these results indicate that PaPaGei-S is relatively robust to skin tone variations, and that additional future work is needed to make it better darker skin tones.
@@ Appendix G Effect of Skin Tone
Appendix H Extended related work
@@
Table 17: Comparison of large-scale PPG studies. * indicates partial availability. The participants and hours indicate pre-training data.
Study #Participants (#Hours) #Devices (Types) Open Data Open Weights Open Code #Tasks (#Datasets)
Abbaspourazad et al. (2023) 141,207 ( [^3] K) 1 (Smartwatch) $>$46 (1)
Ding et al. (2024) 28,539 (300K) 5-6 (ICU, Smartwatch) * 4 (7)
Yun et al. (2024) 170,714 (Varying) [^7] https://biobank.ndph.ox.ac.uk/crystal/field.cgi?id=4205 1 (Finger) ✓* 2 (4)
PaPaGei (Ours) 13,517 (57K) 7 (ICU, Smartwatch, Finger, ”Phone Ox.”) 20 (10)
However, despite the popularity of SSL, there are no widely used FMs for PPG data. While (Abbaspourazad et al., 2023) showcased the potential of foundation models for physiological signals, it was based on a single proprietary dataset and device (Apple Watch) while the models were not released, limiting its practical use in the research community. Similarly, REGLE’s work (Yun et al., 2024) on the UK Biobank dataset showed that embedding PPG signals can improve genetic discovery and risk prediction outcomes. Although parts of that model and pipeline are public, the dataset is not openly accessible, and the primary goal was not to create a foundation model for PPG but rather to focus on genetics. Another work on the same data showed that PPG embeddings are promising for cardiovascular risk prediction (Weng et al., 2024). SiamQuality (Ding et al., 2024) also trained an unreleased model on 36 million PPG signals using proprietary data. Importantly, most of these works pre-trained on a single-device dataset and did not explore out-of-domain datasets or conduct transfer learning experiments, which are crucial for assessing the true generalizability of foundation models. These studies highlight the potential of PPG-based foundation models but also underscore the need for openly available, pre-trained models that can be widely used and adapted by the research community.
@@
In contrast to these generic approaches, our work takes a domain-specific focus. We curate a large pre-training and evaluation benchmark dedicated exclusively to PPG data. While knowledge gained from generic time series foundation models may transfer to domain-specific tasks like PPG, we expect the performance to be limited compared to a model trained specifically on PPG data. Furthermore, foundation models for ECG (McKeen et al., 2024; Song et al., 2024) or EEG (Yuan et al., 2024b) have shown promise but transferring from one domain-specific model (e.g., ECG) to another (PPG) is likely to be even more challenging, as the underlying signal characteristics can be quite different. For instance, Lai et al. (2023) trained a large-scale 12-lead ECG model for detecting 60 diagnostic terms, while McKeen et al. (2024) developed an open-source ECG FM using 1.6 million 12-lead signals. In brain signal analysis, Yuan et al. (2024b) introduced Brant-2, an EEG and SEEG model supporting tasks like sleep staging and seizure detection. Building on this progress, we adopt a domain-specific approach focused on photoplethysmography (PPG) signals. Building on the increasing interest in modality-specific foundation models, our specialized approach allows us to capture nuances and complexities specific to PPG signals.
@@
Appendix I Extended Discussion
In §5.1, we observed that PaPaGei outperforms baselines in at least 14 out of 20 tasks, with average classification and regression improvements of 4.7%-6.3% and 2.9%-4.9%, respectively. PaPaGei-S performed best for cardiovascular parameters like BP, Hypertension, and Avg. HR, which are closely linked to metrics such as sVRI and IPA (Liang et al., 2018b). Additionally, PaPaGei-P surpassed baselines FMs like Moment and is well-suited for tasks such as Smoking and Arousal.
@@ Appendix I Extended Discussion
To provide future direction regarding the use of PaPaGei, we provide some suggestions. For instance, let’s consider the nuMoM2B dataset which consists of pregnant women. PaPaGei-S obtains an AUROC of 0.78 in pregnancy stage classification and 6.05 is gestation age classification. Compared to the pre-training population with diverse age and gender, the nuMoM2B consists of women generally aged between 20-35. Furthermore, the gestation age readings are collected approximately around the first and third trimester. Given these factors, the target nuMoM2B dataset has many variables contributing toward distribution shift. Therefore, PaPaGei-S can be fine-tuned to address the shift in the following ways: (1) We can align the pre-trained embeddings to the nuMoM2B embedding using unsupervised or semi-supervised domain adaptation. (2) Domain Generalization is also an option during the training phase to improve generalization robustness. (3) Newer methods such as LoRA can provide another way to quickly fine-tune. (4) Importantly, given that more women are present in the first visit compared to the third visit, we can optimize different metrics to improve accuracy under the imbalance. For example, AUPRC can be optimized instead of AUROC. Fairness of classification across genders can also be considered during training. Exploring these avenues to further enhance the performance and applicability of PaPaGei is a promising direction for future studies. Moreover, future work may benefit from exploring PPG specific augmentations such as GAN-based approaches (Kiyasseh et al., 2020); and systematically evaluating different augmentations to provide insights into useful PPG augmentations.